BSM-South is due to start up again shortly, as Peter Nelson returns from vacation.
W30 is running, but Bill Goff is in the process of replacing the control computer. This will take a while to complete. In the meantime, he is running the system with the old computer. After a frontal passage, the weather is improving and observations should start shortly.
BSM_Berry has been down for the past week, as we installed new software and hardware. I hope to start observing tonight; the weather is predicted to be clear/partly cloudy most of the week, good enough for more V339 Del observations.
SRO is waking up. John hopes to get down to Sonoita later this week to test things and restart the system.
OC61 is down temporarily. They had high winds (140kph) about a week ago that damaged the dome shutter mechanism. Nigel is rigging up scaffolding to do repairs to the motor shaft. With luck, this should be a short outage.
We hope to get TMO61 and Bareket up and running shortly. BSM_Hamren and BSM_NM will be the next two telescopes that become operational.